OpenStudy (anonymous):

Suppose there are 15 jelly beans in a box (2 red, 3 blue, 4 white, and 6 green). A jelly bean is selected at random. What is the probability that the jelly bean is neither red nor blue?

Nnesha (nnesha):

total : 15 take out the red and blue how many jellybeans would you have left ?

OpenStudy (anonymous):

10

Nnesha (nnesha):

\[\frac{ 15-(\color{ReD}{2}+\color{blue}{3}) }{ 15 }\] yes right!

OpenStudy (anonymous):

so would it be 10/15 = 2/3 in lowest terms?

Nnesha (nnesha):

yes right!

OpenStudy (anonymous):

thank you! very helpful

Nnesha (nnesha):

np :=)